# High-dimensional Laplace asymptotics up to the concentration threshold

### v1

@@ Abstract
$I(\lambda):=\int_{\mathbb{R}^{d}}g(x)e^{-\lambda f(x)}\mathrm{d}x,$
in the regime where $d$ and $\lambda$ are both large. Until now, rigorous bounds for Laplace expansions in growing dimension have been restricted to the “Gaussian-approximation” regime, known to hold when $d^{2}/\lambda\to 0$. This excludes many practically relevant regimes, including those arising in physics and modern high-dimensional statistics, which operate beyond this threshold while still satisfying the concentration condition $d/\lambda\to 0$. Here, we close this gap. We develop an explicit asymptotic expansion for $\log I(\lambda)$ with quantitative remainder bounds that remain valid throughout this intermediate region, arbitrarily close to the concentration threshold $d/\lambda\to 0$.
@@
$\log I(\lambda)=\sum_{k=1}^{L-1}b_{k}(f,g)\lambda^{-k}+\mathcal{O}(d^{L+1}/\lambda^{L}),\quad d^{L+1}/\lambda^{L}\to 0,$
@@
The proof is constructive and proceeds via explicit polynomial changes of variables that iteratively “quadratize” the exponent while controlling Jacobian effects, thereby avoiding heavy Gaussian concentration machinery. We illustrate the expansion on two representative examples.
@@ 1 Introduction
$I(\lambda):=\int_{\mathbb{R}^{d}}g(x)e^{-\lambda f(x)}\mathrm{d}x$ (1.1)
@@
In the classical regime where the dimension $d$ is fixed and $\lambda\to\infty$, Laplace’s method and its higher-order refinements are well developed [4, Section 8.3], [44, Section 5, Chapter IX]. At the opposite extreme, there is also an infinite-dimensional theory (e.g. on Wiener space) in which $\lambda^{-1}$ plays the role of a small-noise parameter [2]. What remains comparatively less understood is the intermediate regime in which the dimension grows with the large parameter. This “growing-dd” regime is now very common in modern statistics, where Laplace-type integrals (or ratios thereof) appear as marginal likelihoods and posterior expectations in models whose parameter dimension increases with sample size. See Section 1.2 for more details. Likewise, the absence of rigorously justified “growing-dd” expansions created a significant void in areas such as statistical physics, Euclidean quantum field theories (QFT), and chemistry, where formal expansions have been in use for a long time without adequate justification, see Section 1.1.
@@
Recent work by the second author establishes a high-dimensional Laplace expansion of $I(\lambda)$, with explicit remainder control, under the scaling $d^{2}/\lambda\to 0$ [22]. The proof centers around an approximation of $f$ in the exponent by its second-order Taylor expansion. Thus the expansion of $I(\lambda)$ is closely tied to the problem of approximating concentrating densities, of the form $\pi\propto e^{-\lambda f}$, by Gaussian distributions. In fact, the condition $d^{2}\ll\lambda$ arises in numerous contexts as a threshold for Gaussian approximation accuracy. In the context of Gaussian approximation to posteriors in Bayesian inference, $d^{2}\ll\lambda$ appears in the works [21, 20, 17, 23]. The condition $d^{2}\ll\lambda$ also arises in proofs of high-dimensional Central Limit Theorems (CLTs); see [32, 46, 19]. In the CLT, $\lambda$ (usually denoted $n$) is the given number of i.i.d. $d$-dimensional random vectors, whose average is of interest. In fact, the aforementioned works [22, 21, 32] establish lower bounds as well, proving $d^{2}/\lambda\ll 1$ is necessary for the Gaussian approximations to be accurate. Although some works have in fact proved the CLT under much weaker conditions on $d$ relative to $\lambda$, the Gaussian approximation holds in a much weaker sense. See Section 3.1 and Remark 2 of [8] for an overview of this line of work.
@@
Thus to summarize, the regime $d^{2}/\lambda\to 0$ arises as a critical threshold in numerous results on approximating $d$-dimensional densities or integrals involving $e^{-\lambda f}$. Another critical threshold is the regime $d/\lambda\to 0$, which is generically required for distributions $\propto e^{-\lambda f}$ to concentrate near the minimizer of $f$; see e.g. [12, Corollary 2.1] and [38, Section 4.1]. Clearly, if there is no concentration around the minimizer of $f$, one should not expect an expansion to hold whose terms are determined solely by the derivatives of $g$ and $f$ at this single point.
@@
The series approximates the logarithm of the integral $I(\lambda)$, and as we explain below, this is precisely what allows us to push $d$ above the $\sqrt{\lambda}$ barrier. Let $x_{\star}$ be the minimizer of $f$. Assuming bounded operator norms of derivatives of $f$ and $g$ near $x_{\star}$ through orders $2L+2$ and $2L$, respectively, and mild global growth of $f$ away from $x_{\star}$, we show that
@@
for some coefficients $b_{k}(f,g)$ that satisfy $b_{k}(f,g)=\mathcal{O}(d^{k+1})$. If $d\lesssim\log\lambda$, then additional factors of $\log\lambda$ are present in the remainder. The sum with respect to $k$ is omitted if $L=1$, and we assumed without loss of generality that $g=1$ at $x_{\star}$. We prove that the coefficients $b_{k}$ can be derived from formal (and well-known) cumulant expansions.
@@
A central significance of our result is that it advances the modern asymptotic analysis program of developing expansions of ubiquitous Laplace type integrals that remain accurate in increasingly high-dimensional regimes. The best known results to date work under the assumption $d^{2}/\lambda\to 0$. In contrast, we extend the range of dimensions for which one can make precise asymptotic statements with explicit formulas until the very limit, because beyond $d/\lambda\to 0$ there is no concentration any longer. In this sense, our work brings the program initiated by Laplace in the 1700s to its logical conclusion, by providing explicit asymptotics (with remainder control) up to the threshold where Laplace-type concentration breaks down. We show that a simple analytic approximation remains valid whenever $d^{L+1}/\lambda^{L}\to 0$ for any $L\geq 1$, even in the genuinely intermediate region where $d^{2}/\lambda\not\to 0$ (indeed, where $d^{2}/\lambda$ may diverge), by identifying an explicit exponential correction at the level of the log-integral.
@@
The first application is in physics, including statistical physics and QFT, where Laplace-type integrals encode quantities of central importance. In the high-dimensional, many-degrees-of-freedom regimes relevant to these fields, such quantities are often evaluated via formal Laplace expansions, typically without remainder bounds that remain valid in the joint limit of a large parameter and growing dimension; see Section 1.1 for details. This theoretical gap, which dates back at least to the Darwin–Fowler steepest-descent approach in 1922 [9], has long limited the rigor of many computations. Our results fill the century-old gap by placing these Laplace calculations on firm mathematical footing for a broad class of finite-dimensional large-system models.
Another area where our results are of significant importance is statistics. Here, $\lambda$ plays the role of sample size. Our results address the need of modern statistics to develop rigorously-justified Laplace asymptotics that remain accurate in increasingly high-dimensional regimes. Indeed, high-dimensional Laplace integrals are ubiquitous in statistics, arising as marginal likelihoods, model evidences, or other posterior expectations. Complicating matters is the fact that these integrals must often be evaluated repeatedly for different parameter choices, and then optimized over. Laplace approximations such as the Bayesian Information Criterion are popular and extremely convenient because they produce explicit approximations rather than black-box numerical estimates. Yet until now, theoretical backing for these approximations was limited to the $d^{2}/\lambda\to 0$ regime. Our results push these guarantees into substantially higher-dimensional regimes, arbitrarily close to $d/\lambda\to 0$. See Section 1.2 for more details and references.
@@
Finally, our result may also be of interest in the theory of cumulants, since it solves the problem of bounding the remainder in cumulant expansions. Obtaining a remainder bound using cumulant theory (and the closely related theory of Gaussian chaos [15, 31, 30]) directly is deeply nontrivial. We have avoided this problem by finding an alternative route. See Section 1.3 for more details.
@@ 1.1 Application in physics
plays a central role. Here $\beta$ is inversely proportional to the temperature, $\mathcal{H}$ is the Hamiltonian of the model, and $\int\dots\mathcal{D}\phi$ denotes functional (infinite-dimensional) integration over all allowed configurations $\phi$ of the model [6, Sections 2.1], [16, Sections 2.1], [47, Introduction to Chapter 2]. The normalized logarithm of the partition function is known as free energy: $\mathcal{F}=-\beta^{-1}\log Z$. The functional integral in (1.3) is typically interpreted as the limit of the appropriate finite dimensional integrals [16, Sections 2.3]. After discretization one obtains an ordinary (finite-dimensional) integral over $d$ degrees of freedom,
@@
where $\lambda$ is a large prefactor coming from the problem [16, Sections 2.3]. If the model is discrete to begin with (i.e., it has finitely many degrees of freedom $d$), one gets (1.4) right away [16, Sections 2.1 and 2.2] and [6, Sections 1.1–1.3].
@@
In physics texts, this step is typically justified by a combination of (i) the presence of a large parameter (e.g. volume), and (ii) an a posteriori self-consistency check that the discarded terms are “small” in the regime of interest; see again [16, Sec. 2.3] for an example of this approach.
@@
Taking logs yields a formal cumulant expansion [26, Section 3.10], [11, Sections 3.1–3.6]:
@@
where $\mathrm{cum}$ denotes the corresponding cumulant. In field-theoretic language, the terms in (1.8) are precisely the loop corrections to mean field: Wick expansion of the Gaussian moments [16, Sections 5.1], [47, Section 1.1] produces a sum over Feynman diagrams, and the cumulant picks out the connected diagrams, which correct the free energy $\mathcal{F}_{d}(\lambda)$ order-by-order in $\lambda^{-1}$ [6, Sections 9.1], [47, Section 1.2 and Chapter 7], [11, Sections 3.13].
Physics literature makes extensive use of truncations of (1.8) (or equivalent loop expansions). However, derivations are usually done without rigorous error control [16, Sections 2.1 and 2.2] and [6, Sections 1.1–1.3]. For finite $d$, the error control follows from classical asymptotic analysis [4, 44]. For growing $d$, such expansions have never been justified except for some specific cases.
@@ 1.2 Application in statistics
Laplace-type integrals $\int_{\mathbb{R}^{d}}g(x)e^{-nf(x)}\mathrm{d}x$ are omnipresent in statistics, with $n=\lambda$ typically playing the role of sample size, and $f$ depending weakly on $n$ [40]. In Bayesian inference, $e^{-nf(x)}=P(\mathrm{data}\mid x)P_{0}(x)$. The probability distribution $P(x\mid\mathrm{data})=e^{-nf(x)}/\int e^{-nf(x^{\prime})}\mathrm{d}x^{\prime}$ is the posterior probability distribution of an unknown parameter $x$ given $\mathrm{data}$, consisting of $n$ independent data points. Once the posterior has been specified, one is interested in its summary statistics, such as the posterior mean and variance-covariance matrix, which can be expressed as ratios of Laplace-type integrals $\int_{\mathbb{R}^{d}}g(x)e^{-nf(x)}\mathrm{d}x/\int_{\mathbb{R}^{d}}e^{-nf(x)}\mathrm{d}x$ [40, 41].
Another target for computation is the denominator $\int e^{-nf(x)}\mathrm{d}x$ itself. This quantity is known as the model evidence [18]. Different modeling choices for $P(\mathrm{data}\mid x)$ (how the data is generated from the unknown parameter $x$), as well as for $P_{0}(x)$ (the prior probability distribution of $x$), lead to different values of the model evidence. Thus a model can be chosen among several, or among a continuum, by maximizing $\mathcal{M}\mapsto\int e^{-nf(x)}\mathrm{d}x$. This is known as Bayesian model selection [18, 43][1, Chapter 5].
In frequentist inference, Laplace-type integrals arise as marginal likelihoods [34]. Specifically, suppose we are interested in estimating a parameter $\theta\in\mathbb{R}^{k}$, but we have a nuisance parameter $x\in\mathbb{R}^{d}$, which in turn is drawn from a distribution $g$. Then the likelihood of the data takes the form
$L(\theta)=P(\mathrm{data}\mid\theta)=\int_{\mathbb{R}^{d}}P(\mathrm{data}\mid x,\theta)q(x)\mathrm{d}x=\int g(x)e^{-nf(x,\theta)}\mathrm{d}x$
[36]. Here, the first two equalities are by definition, while the integrand in the rightmost integral expresses the mathematical structure of $P(\mathrm{data}\mid x,\theta)$. One is now interested in the marginal maximum likelihood estimator $\hat{\theta}=\arg\max_{\theta}L(\theta)=\arg\max_{\theta}\int g(x)e^{-nf(x,\theta)}\mathrm{d}x$.
To summarize, we are interested either in approximations of $\int g(x)e^{-nf(x)}\mathrm{d}x$ for a particular $g$ and $f$, or in maximizing a function $\theta\mapsto I_{n}(\theta)=\int_{\mathbb{R}^{d}}g(x)e^{-nf(x,\theta)}\mathrm{d}x$ (Bayesian model selection can be classified as the latter task). These tasks become particularly demanding in the growing-$d$ regimes that now arise routinely in applications. For example, [14] tackles the above Laplace optimization problem in the context of deep learning, where the integration is over the space of hundreds of parameters of a neural network.

### v2

@@ Abstract
$I(\lambda):=\left(\frac{\lambda}{2\pi}\right)^{d/2}\int_{\mathbb{R}^{d}}g(x)e^{-\lambda f(x)}\mathrm{d}x,$
in the regime where $d$ and $\lambda$ are both large. Until now, rigorous bounds for the Laplace expansion in growing dimension have been restricted to the “Gaussian-approximation” regime, known to hold when $d^{2}/\lambda\to 0$. This excludes many practically relevant regimes, including those arising in physics and modern high-dimensional statistics, which operate beyond this threshold while still satisfying the concentration condition $d/\lambda\to 0$. Here, we close this gap. We develop an explicit asymptotic expansion for $\log I(\lambda)$ with quantitative remainder bounds that remain valid throughout this intermediate region, arbitrarily close to the concentration threshold $d/\lambda\to 0$.
@@
$\log I(\lambda)=\sum_{k=1}^{L-1}b_{k}(f,g)\lambda^{-k}+\mathcal{O}(d^{L+1}/\lambda^{L}),\qquad d^{L+1}/\lambda^{L}\to 0,$ (0.1)
@@
In addition, we study the problem of computing expectations against, and sampling from, concentrating Laplace-type probability densities $\pi(x)\propto e^{-\lambda f(x)}$. For computing expectations of smooth observables $g$, we propose an approximation based on (LABEL:I-abstract). For sampling, we construct a family of push-forward densities $\hat{\pi}_{L}:=(x_{L})_{\#}\mathcal{N}(0,\lambda^{-1}I_{d})$, $L=1,2,3\dots$ approximating $\pi$ with accuracy $\operatorname{TV}(\pi,\hat{\pi}_{L})\lesssim d^{L+1}/\lambda^{L}$. Here, the maps $x_{L}$ are explicit polynomials. By taking $L$ large enough, here too, we can take $d$ arbitrarily close to the concentration threshold $d=o(\lambda)$.
@@ 1 Introduction
$I(\lambda):=\left(\frac{\lambda}{2\pi}\right)^{d/2}\int_{\mathbb{R}^{d}}g(x)e^{-\lambda f(x)}\mathrm{d}x$ (1.1)
@@
In the classical regime where the dimension $d$ is fixed and $\lambda\to\infty$, Laplace’s method and its higher-order refinements are well developed [5, Section 8.3], [51, Section 5, Chapter IX]. At the opposite extreme, there is also an infinite-dimensional theory (e.g. on Wiener space) in which $\lambda^{-1}$ plays the role of a small-noise parameter [2]. What remains comparatively less understood is the intermediate regime in which the dimension grows with the large parameter. This “growing-dd” regime is now very common in modern statistics, where Laplace-type integrals (or ratios thereof) appear as marginal likelihoods and posterior expectations in models whose parameter dimension increases with sample size. See Section 1.2 for more details. Likewise, the absence of rigorously justified “growing-dd” expansions created a significant void in areas such as statistical physics, Euclidean quantum field theories (QFT), and chemistry, where formal expansions have been in use for a long time without adequate justification, see Section 1.1.
@@
Recent work by the second author establishes a high-dimensional Laplace expansion (LE) of $I(\lambda)$, with explicit remainder control, under the scaling $d^{2}/\lambda\to 0$ [28]. The proof centers around an approximation of $f$ in the exponent by its second-order Taylor expansion. Thus the expansion of $I(\lambda)$ is closely tied to the problem of approximating concentrating densities, of the form $\pi\propto e^{-\lambda f}$, by Gaussian distributions. In fact, the condition $d^{2}\ll\lambda$ arises in numerous contexts as a threshold for Gaussian approximation accuracy. In the context of Gaussian approximation to posteriors in Bayesian inference, $d^{2}\ll\lambda$ appears in the works [27, 26, 22, 30]. The condition $d^{2}\ll\lambda$ also arises in proofs of high-dimensional Central Limit Theorems (CLTs); see [40, 53, 24]. In the CLT, $\lambda$ (usually denoted $n$) is the given number of i.i.d. $d$-dimensional random vectors, whose average is of interest. In fact, the aforementioned works [28, 27, 40] establish lower bounds as well, proving $d^{2}/\lambda\ll 1$ is necessary for the Gaussian approximations to be accurate. Although some works have in fact proved the CLT under much weaker conditions on $d$ relative to $\lambda$, the Gaussian approximation holds in a much weaker sense. See Section 3.1 and Remark 2 of [9] for an overview of this line of work.
@@
Thus to summarize, the regime $d^{2}/\lambda\to 0$ arises as a critical threshold in numerous results on approximating $d$-dimensional densities or integrals involving $e^{-\lambda f}$. Another critical threshold is the regime $d/\lambda\to 0$, which is generically required for distributions $\propto e^{-\lambda f}$ to concentrate near the minimizer of $f$; see e.g. [17, Corollary 2.1] and [45, Section 4.1]. Clearly, if there is no concentration around the minimizer of $f$, one should not expect an expansion to hold whose terms are determined solely by the derivatives of $g$ and $f$ at this single point.
@@
The series approximates the logarithm of the integral $I(\lambda)$, and as we explain below, this is precisely what allows us to push $d$ above the $\sqrt{\lambda}$ barrier. Let $x_{\star}$ be the global minimizer of $f$ and assume without loss of generality that $\nabla^{2}f(x_{\star})=I_{d}$. Assuming bounded operator norms of derivatives of $f$ and $g$ near $x_{\star}$ through orders $2L+2$ and $2L$, respectively, and mild global growth conditions on $f$ and $g$, we show that
@@
for some coefficients $b_{k}(f,g)$ that satisfy $b_{k}(f,g)=\mathcal{O}(d^{k+1})$. If $d\lesssim\log\lambda$, then additional factors of $\log\lambda$ are present in the remainder. The sum with respect to $k$ is omitted if $L=1$, and we assumed without loss of generality that $g=1$ at $x_{\star}$. The coefficients $b_{k}$ are already well-known, and can be derived from formal cumulant expansions [34, Section 3.10].
In fact, more broadly, our contribution is a powerful technique to tackle Laplace integrals. We use this technique not only to prove (1.2), but also to solve a related problem of approximating a Laplace-type probability density, of the form $\pi(x)\propto e^{-\lambda f(x)}$. Namely, we construct a transformation $x_{L}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\hat{\pi}_{L}:=(x_{L})_{\#}\mathcal{N}(0,\lambda^{-1}I_{d})$ approximates $\pi$:
$\mathrm{TV}(\pi,\hat{\pi}_{L})\lesssim d^{L+1}/\lambda^{L}.$ (1.3)
Here, $\mathcal{N}(0,\lambda^{-1}I_{d})$ is the Gaussian distribution with zero mean and covariance matrix $\lambda^{-1}I_{d}$, while TV stands for the total variation distance. The subscript #\# in the definition of $\hat{\pi}_{L}$ denotes the push-forward. This means $\hat{\pi}_{L}$ is the law of the random variable $x_{L}(Z)$ when $Z$ is distributed as $\mathcal{N}(0,\lambda^{-1}I_{d})$. The reason $\hat{\pi}_{L}$ is so useful is that it is easy to sample from, precisely due to this push-forward construction. Cheaply generating approximate samples from an untractable density $\pi\propto e^{-\lambda f}$ is an important problem in Bayesian statistics. See Section 1.2 for more details.
@@
A central significance of our result is that it advances the modern asymptotic analysis program of developing expansions of ubiquitous Laplace type integrals that remain accurate in increasingly high-dimensional regimes. The best known results to date work under the assumption $d^{2}/\lambda\to 0$. In contrast, we extend the range of dimensions for which one can make precise asymptotic statements with explicit formulas until the very limit, because beyond $d/\lambda\to 0$ there is no concentration any longer. In this sense, our work essentially completes the classical Laplace program for concentrating finite-dimensional integrals under natural smoothness and growth assumptions, by proving remainder bounds on the high-dimensional LE up to the concentration threshold. We show that a constructive analytic approximation remains valid whenever $d^{L+1}/\lambda^{L}\to 0$ for any $L\geq 1$, even in the genuinely intermediate region where $d^{2}/\lambda\not\to 0$ (indeed, where $d^{2}/\lambda$ may diverge), by identifying an explicit exponential correction at the level of the log-integral.
@@
The first application is in physics, including statistical physics and QFT, where Laplace-type integrals encode quantities of central importance. In the high-dimensional, many-degrees-of-freedom regimes relevant to these fields, such quantities are often evaluated via formal LEs, typically without appropriate remainder bounds; see Section 1.1 for details. This theoretical gap, which dates back at least to the Darwin–Fowler steepest-descent approach in 1922 [10], has long limited the rigor of many computations. Our results fill the century-old gap by placing these Laplace calculations on firm mathematical footing for a broad class of finite-dimensional large-system models.
Another area where our results are of significant importance is statistics. Here, $\lambda$ plays the role of sample size, and high-dimensional Laplace-type integrals arise ubiquitously as normalizing constants, marginal likelihoods, and posterior expectations. As mentioned above, a central task is to approximate the posterior density $\pi\propto e^{-\lambda f}$: one wants to sample from it, compute expectations of observables against it, and evaluate its normalizing constant for model comparison. Our results address all three. We construct an explicit approximation $\hat{\pi}_{L}$ to $\pi$ from which one can easily sample, we provide closed-form approximations to posterior expectations of smooth observables that avoid Monte Carlo error entirely, and we give rigorous asymptotic expansions of the normalizing constant that generalize the Bayesian Information Criterion (BIC) to higher order. Laplace-based surrogates such as the BIC are popular precisely because they provide explicit analytic approximations rather than black-box numerical estimates. Yet until now, theoretical backing for these approximations was limited to the $d^{2}/\lambda\to 0$ regime. Our results push these guarantees into substantially higher-dimensional regimes, arbitrarily close to the concentration threshold $d/\lambda\to 0$. See Section 1.2 for more details and references.
@@
Finally, our result has interesting connections to the theory of cumulants and to normalizing flows in machine learning. The connection with the former is that we have solved the problem of bounding the remainder in cumulant expansions. Obtaining a high-dimensional remainder bound using cumulant theory (and the closely related theory of Gaussian chaos [20, 39, 38]) directly is deeply nontrivial and has not been done before. We have avoided this problem by finding an alternative route. See Section 1.3 for more details.
Regarding normalizing flows, these are sequences of maps $T_{1},\dots,T_{L}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ with the property that if $Z\sim\mathcal{N}(0,I_{d})$ then $T_{1}\circ\dots\circ T_{L}(Z)$ is approximately distributed according to a target distribution $\pi$ [31, 41]. Typically, $T_{1},\dots,T_{L}$ are constructed using deep neural networks, the parameters of which are found by minimizing a loss [12]. Here, we approximate $\pi$ in a similar fashion, pushing forward a Gaussian distribution through a series of transformations. But the key differences with normalizing flows are that 1) our construction is explicit, not through minimizing some loss, and 2) we prove a rigorous error bound. Essentially the only structural assumption we need to obtain this result is that the target $\pi$ is a concentrating measure.
@@ 1.1 Application in physics
plays a central role. Here $\beta$ is inversely proportional to the temperature, $\mathcal{H}$ is the Hamiltonian of the model, and $\int\dots\mathcal{D}\phi$ denotes functional (infinite-dimensional) integration over all allowed configurations $\phi$ of the model [7, Sections 2.1], [21, Sections 2.1], [54, Introduction to Chapter 2]. The normalized logarithm of the partition function is known as free energy: $\mathcal{F}=-\beta^{-1}\log Z$. The functional integral in (1.4) is typically interpreted as the limit of the appropriate finite dimensional integrals [21, Sections 2.3]. After discretization one obtains an ordinary (finite-dimensional) integral over $d$ degrees of freedom,
@@
where $\lambda$ is a large prefactor coming from the problem [21, Sections 2.3]. If the model is discrete to begin with (i.e., it has finitely many degrees of freedom $d$), one gets (1.5) right away [21, Sections 2.1 and 2.2] and [7, Sections 1.1–1.3].
@@
In physics texts, this step is typically justified by a combination of (i) the presence of a large parameter (e.g. volume), and (ii) an a posteriori self-consistency check that the discarded terms are “small” in the regime of interest; see again [21, Sec. 2.3] for an example of this approach.
@@
Taking logs yields a formal cumulant expansion [34, Section 3.10], [15, Sections 3.1–3.6]:
@@
where $\mathrm{cum}$ denotes the corresponding cumulant. In field-theoretic language, the terms in (1.9) are precisely the loop corrections to mean field: Wick expansion of the Gaussian moments [21, Sections 5.1], [54, Section 1.1] produces a sum over Feynman diagrams, and the cumulant picks out the connected diagrams, which correct the free energy $\mathcal{F}_{d}(\lambda)$ order-by-order in $\lambda^{-1}$ [7, Sections 9.1], [54, Section 1.2 and Chapter 7], [15, Sections 3.13].
Physics literature makes extensive use of truncations of (1.9) (or equivalent loop expansions). However, derivations are usually done without rigorous error control [21, Sections 2.1 and 2.2] and [7, Sections 1.1–1.3]. For finite $d$, the error control follows from classical asymptotic analysis [5, 51]. For growing $d$, such expansions have never been justified except for some specific cases.
@@ 1.2 Application in statistics
For the sake of brevity, we focus on Laplace-type integrals and densities in the context of Bayesian statistics; however, these quantities frequently arise in frequentist statistics as well.
1.2.1 Statistics context and state of the art
Laplace-type integrals $\int_{\mathbb{R}^{d}}g(x)e^{-\lambda f(x)}\mathrm{d}x$ and probability distributions $\pi(x)\propto e^{-\lambda f(x)}$ are omnipresent in statistics. Here, $\lambda$ (more commonly denoted $n$ in statistics) plays the role of sample size, and $f$ depends weakly on $\lambda$ [47]. In Bayesian inference, $\pi(x)=P(x\mid\mathrm{data})$ is the posterior probability distribution of an unknown parameter $x$ given $\mathrm{data}$, consisting of $\lambda=n$ independent data points. Once the posterior has been specified, one is typically interested in (1) computing summary statistics, which take the general form $\mathbb{E}_{X\sim\pi}[g(X)]$, (2) sampling from $\pi$, and (3) computing the normalizing constant $\int e^{-\lambda f(x)}\mathrm{d}x$ itself [16]. Summary statistics distill information in the posterior; samples enable exploration and uncertainty quantification. The normalizing constant, known as the model evidence [23], underpins Bayesian model selection. Here, one finds the best model by optimizing the evidence over candidate models [23, 50][1, Chapter 5]. All three tasks become particularly demanding in the growing-$d$ regime that now arises routinely in applications.
[…]
While approximating $\pi$ by a tractable $\hat{\pi}$ is natural for sampling and for evaluating $\mathbb{E}_{X\sim\pi}[g(X)]$ when $g$ is nonsmooth, one can often do better when $g$ is smooth by working directly with the LE. Indeed, $\mathbb{E}_{X\sim\pi}[g(X)]$ is a ratio of two Laplace-type integrals. Expanding both the numerator and the denominator and then taking the ratio yields an explicit, fully deterministic approximation to $\mathbb{E}_{X\sim\pi}[g(X)]$, avoiding the Monte Carlo error inherent in estimating expectations via samples from $\hat{\pi}$. Despite this advantage, LEs have received comparatively little attention as a tool for computing posterior expectations. Notable exceptions are the fixed-$d$ analyses in [47, 48].